# An Agnostic Biosignature Based on Modeling Panspermia and Terraformation

### v2

@@ Abstract
A fundamental goal of astrobiology is to detect life outside of Earth. This proves to be an exceptional challenge outside of our solar system, where strong assumptions must be made about how life would manifest and interact with its planet. Such assumptions are required because of the lack of a consensus theory of living systems, or an understanding of the possible extent of planetary dynamics. Here we explore a model of life spreading between planetary systems via panspermia and terraformation. Our model shows that as life propagates across the galaxy, correlations emerge between planetary characteristics and location, and can function as a population-scale agnostic biosignature. This biosignature is agnostic because it is independent of strong assumptions about any particular instantiation of life or planetary characteristic—by focusing on a specific hypothesis of what life may do, rather than what life may be. By clustering planets based on their observed characteristics, and examining the spatial extent of these clusters, we demonstrate (and evaluate) a way to prioritize specific planets for further observation—based on their potential for containing life. We consider obstacles that must be overcome to practically implement our approach, including identifying specific ways in which better understanding astrophysical and planetary processes would improve our ability to detect life. Finally, we consider how this model leads us to think in novel ways about scales of life and planetary replication.
Panspermia — Terraformation — Astronomical Simulation
1 Introduction
@@
It is difficult to attribute, with certainty, observable features of exoplanets to extraterrestrial life (Moore et al., 2017; Tasker et al., 2017; Green et al., 2021; Cockell, 2022; Lenardic et al., 2022, 2023; Smith & Mathis, 2023). This is exemplified by the insufficient (and often ill-defined) working definitions of life that are used to interpret observational data, and by the growing number of false positives for traditional biosignatures (Cleland & Chyba, 2002; Benner, 2010; Cleland, 2012; Mix, 2015; Schwieterman et al., 2016; Bich & Green, 2018; Harman & Domagal-Goldman, 2018; Schwieterman et al., 2018; Mariscal & Doolittle, 2020; Janin, 2021; Vickers et al., 2023).
@@
One way in which the community aims to overcome the problems with defining life is to develop “agnostic” biosignatures—detecting signs of life that are not particular to Earth-life or any other hypothetical instances of life (Marshall et al., 2021; Smith et al., 2021). Yet, agnostic biosignature proposals are sometimes built on restricted concepts of habitability (e.g., requiring rocky, watery planets), or simple anomaly detection (i.e., without using a working definition of life) (Kinney & Kempes, 2022; Cleaves et al., 2023).
@@
A proposed remedy for the issues surrounding traditional biosignatures has been to use so-called “statistical” biosignatures—relying on integrating multiple lines of evidence to increase confidence in a discovery, or generating ensembles of data to better constrain the probabilities and (in Bayesian speak). (Lin & Loeb, 2015; Catling et al., 2018; Walker et al., 2018; Affholder et al., 2021; Bixel & Apai, 2021). However, such statistical biosignatures often rely on assumptions about the prior probability of abiogenesis, or on the existence of unambiguous biosignatures for single planets. These approaches are often trying to solve more specific issues, like estimating the frequency of Earth-like life, the frequency of planets originating from panspermia compared to abiogenesis, or finding evidence of a specific metabolic process (Lin & Loeb, 2015; Walker et al., 2018; Affholder et al., 2021; Checlair et al., 2021; Kovacevic, 2022).
@@
Here we ask, can we detect the presence of life if we postulate that life is spreading between and terraforming planets?111Here, “life spreading” refers to interstellar panspermia, and “terraforming planets” refers to modifying observable characteristics of planets. This is an astrobiological “hinge proposition” of the kind described by Kinney & Kempes as necessary to make sure “that the door of astrobiology can turn properly” under conditions of deep uncertainty (Kinney & Kempes, 2022). It is also somewhat the converse of a question posed in other panspermia related work: Can the prevalence of panspermia be constrained when assuming we have a way to detect life? (Lin & Loeb, 2015; Balbi & Grimaldi, 2020; Grimaldi et al., 2021; Lingam et al., 2021; Kovacevic, 2022).
@@
In fact, our postulates (of terraformation and panspermia) are less peculiar than they might at first seem: while existing work doesn’t always explicitly refer to terraformation, it must be assumed—in the sense that the mere presence of life stable at geological timescales would create environmental feedback with a planet, and if that life is to be detectable it must modify a planet’s observables. Such planetary modification by life is a well documented phenomena—e.g., the rise of during the great oxygenation event (Lynas et al., 2021), or the rise of from human industrial activity (Olejarz et al., 2021)).
@@
The feasibility of interstellar lithopanspermia (non-intelligent exchange of material via rocks) has been discussed at length in other work (see e.g. Carroll-Nellenback et al. (2019); Grimaldi et al. (2021); Gobat (2021); Totani (2023) and references therein). Crucially, it appears plausible, although likelihoods and rates vary considerably depending on assumptions on timing, amount of material ejected, organismal hardiness, and capture rates, among other features. Arguments have likewise been made for the feasibility of directed interstellar panspermia by intelligent life (Wright et al., 2014; Newman & Sagan, 1981)Ultimately, our postulates of panspermia and terraformation are merely well understood hallmarks of life (proliferation via replication, and adaptation with bi-directional environmental feedback), escalated to the planetary scale, and executed on an interstellar scale.
@@
We use a simulated model to show that statistical correlations between the spatial distribution of planets around different host stars, and their observable characteristics would itself be evidence of life, without the need for a separate biosignature that could reliably detect life on any given planet in isolation. The agnosticism of this biosignature is inseparable from its emergence at the scale of a population of planets—singleton planetary anomalies might be explained away by unknown geochemical processes, or targeted simply because they are anomalous (without a clear hypothesis of why they should be explained by life). Hypothesizing that life spreads via panspermia and terraformation allows us to search for biosignatures while forgoing any strong assumptions about not only the peculiarities of life (e.g., its metabolism) and planetary habitability (e.g., requiring surface liquid water) (Gobat, 2021), but even the potential breadth of structure and chemical complexity underpinning living systems (Solé & Munteanu, 2004; Kim et al., 2019; Bartlett et al., 2022; Wong et al., 2023).
@@
We first describe our model (Sec. 2.1) and approach for statistically identifying the presence of terraformed planets (Sec. 2.2). We show that evolving our model in time can increase the correlation between planets’ compositions and positions (Sec. 3.1). To identify specific planets likely to have been terraformed, we cluster planets by observable characteristics, then select clusters which are spatially localized and cause a decrease in correlation when removed (Secs. 2.2.3, 2.2.4 ). We evaluate these clusters by calculating how well they correctly classify terraformed planets as being terraformed, and non-terraformed planets as being non-terraformed (Sec. 3.2). Finally, we discuss how our results might change due to theoretical and observational constraints, identify specific ways in which better understanding astrophysical and planetary processes could improve life detection, and speculate on the concept of life at the scale of populations of planets.
2 Methods
2.1 Modeling Panspermia and Terraformation
2.1.1 Summary
@@
We created an agent-based model to simulate life222In this work, we use life specifically to refer to an agent (used strictly in the “entity” sense of the word agent) traveling from a parent terraformed planet to a target non-terraformed planet, and whose arrival at the target planet triggers terraformation. spreading between planets orbiting different host stars. The model is initialized with 1000 planets (1 of which is terraformed/panspermia-capable), fixed in a frame of reference, and uniformly randomly distributed in a continuous 3D volume. All planets are assumed to be in different star systems (i.e., we assume one planet per star). “Planets” have compositions representing the observable characteristics of the planet, and “life” agents have compositions representing the observable characteristics of the planet which the life originates from. During simulation, terraformed planets 333We specifically avoid saying living planets, because our methodology cannot distinguish between living planets incapable of panspermia and non-living planets. emit life towards a target non-terraformed planet at a constant velocity 444This approach is most similar to directed panspermia by intelligent life, but could also be envisioned as undirected lithopanspermia, where we refrain from burdening the simulations with agents that won’t ever be able to interact with planets under the rules of our model.. The target planets are selected based on proximity in position and in composition (see Sec. 2.1.4). When life arrives at a planet, the planet’s composition is modified (terraformed) based on both the initial composition of the planet and the composition of the incoming life. Life originating from that terraformed planet will then reflect its new composition. We assume that the time required for terraformation is negligible compared to the time required to travel between planets. Terraformed planets emit life at a fixed rate, provided suitable target planets exist (see Sec. 2.1.4). The simulation ends when all planets are terraformed, or no suitable target planets remain (Fig. 1B).
2.1.2 Compositions and Terraformation
@@
Both “life” and “planets” have compositions represented by vectors of 10 real numbers (each element ). In our model, these compositions are abstractions, used as a means to compare the compatibility of any given life and planet. While we might in the abstract think of one vector representing the observable characteristics of a planet, this same vector would not represent the same observable characteristics of life compatible with that planet. For life, it instead represents the kind of planetary observable characteristics which are compatible with that life. For example, perhaps a planet’s composition vector corresponds to an atmosphere composed of 95% and 5% . The same composition vector in the context of life is used to refer to a kind of life which could be compatible with a planet with an atmospheric composition of 95% and 5% . This does not mean that life on such a planet is literally composed of 95% and 5% . This captures the idea that if a life and a planet have identical composition vectors, then that life will surely be able to terraform the planet555This is an extreme example where technically life wouldn’t even need to terraform the planet, because it implies that the planet’s observable characteristics already perfectly match the planet from where the life originated.. Correspondingly, a life’s composition is identical to the planet from which it emits. Depending on the simulation parameters, initial compositions of each planet are randomly drawn from a uniform distribution. This is a loose approximation of the unknowns surrounding the distribution of diversity of planets between planetary systems (Tasker et al., 2020; Pacetti et al., 2022).
@@
Planet compositions are modified by the process of terraformation, where a mixing function determines how to generate the terraformed composition based on the non-terraformed planet composition and the incoming life’s composition. We apply a mixing function inspired by horizontal gene transfer, but other mixing functions could also be used. The most important feature we aim to capture is the bi-directional feedback between the terraforming life, and the terraformed planet—such that future life emitted from the terraformed planet will inherit the post-terraformation characteristics of that planet. We keep randomly chosen element from the planet’s pre-terraformed composition, and the rest of the terraformed composition is inherited from the terraforming life (Fig. 1C).
2.1.3 Model Stepping
@@
Each step of the model increments the time by . Time units are arbitrary, and only impact the distance that life travels in a step, . Here, , and , meaning that for every 10 model steps, a planet can emit 1 new life, and existing life can travel 20 units.
2.1.4 Target Planet Selection
@@
The destination selection aims to balance the need of identifying a target planet that is both compositionally similar, and physically near, the parent planet. This is based on the assumption that life directing panspermia would want to minimize travel time and terraformation energy costs. In the context of lithopanspermia, this could be interpreted as an assumption that beyond a certain positional distance capture becomes vanishingly unlikely, and beyond a certain compositional distance compatibility becomes vanishingly unlikely.
@@
The algorithm we used for determining a life’s target planet has a positional distance threshold, . Among all non-terraformed planets, , under that positional distance threshold from the life’s origin planet, it selects the planet nearest in compositional space, . Here, , is shorthand for all non-terraformed planets within of life’s origin planet, and is the planet at minimum compositional distance to the life’s origin planet (Fig. 1A). The distances calculated are Euclidean. A range of positional thresholds, , corresponding to and of the maximum distance possible between any two points in space were chosen for in order to span the transition between a model where terraformation of every planet is impossible and every planet is possible, based on the average distance between planets.
2.2 Identifying the Presence of Terraformed Planets
2.2.1 Summary
@@
We hypothesize that the process of panspermia and terraformation in our model will lead to a population of planets with anomalously high positive correlations between their spatial locations and compositions, compared to random permutations of these planets’ compositions. We quantify this using the Mantel test—a statistical test common in ecological science (Sec. 2.2.2). By clustering on the planet compositions, we can begin to pin down clusters of planets driving these correlations (Sec. 2.2.3). From these initial clusters, we select those localized in space (via their interquartile range, IQR), because we hypothesize that life would not only change the distribution of planetary compositions, but would also do so in a relatively compact portion of the galaxy. We further select clusters which, when removed, cause a decrease in the Mantel coefficient of the residual space of planets (Sec. 2.2.4). Finally, we attempt to evaluate how well our clusters reflect the presence of truly terraformed planets (Sec. 2.2.5).
2.2.2 Mantel Test

### v3

# An Agnostic Biosignature Based on Modeling Panspermia and Terraforming

@@ Abstract
The search for a second instance of life is one of the greatest problems of modern science. Outside of creating an artificial origin of life on Earth, the primary targets for the search for life are planets inside or outside the solar system. Realistically, there are just a few locations to search for alien life within the solar system. Outside the solar system, opportunities are nearly unlimited, but there’s a catch: it is difficult to attribute, with certainty, features of exoplanets to extraterrestrial life. Simple spectral biosignatures are susceptible to false positives; technosignatures reduce this susceptibility at the expense of strong assumptions about potential underlying life and its technologies. We have developed an agnostic approach to exoplanet life detection that overcomes these limitations by using properties that emerge on the scale of groups of planets, without the need for a “smoking-gun” single-planet level biosignature. We use an agent-based model to show that if life can spread between star systems, and affect the observable properties of a planet, then a robust signature of life (with very few false positives) can emerge, defined by correlations between planet characteristics and their locations. By clustering planets based only on their observed characteristics, and retaining clusters localized in space, we demonstrate (and evaluate) a way to prioritize specific planets for further observation, based on their potential for containing life. We consider obstacles that must be overcome to practically implement our approach, including identifying specific ways in which better understanding astrophysical and planetary processes would improve our ability to detect life.
Panspermia — Terraforming — Astronomical Simulation
I Introduction
@@
It is difficult to attribute, with certainty, observable features of exoplanets to extraterrestrial life (Moore et al., 2017; Tasker et al., 2017; Green et al., 2021; Cockell, 2022; Lenardic et al., 2022, 2023; Smith and Mathis, 2023). This is exemplified by the insufficient (and often ill-defined) working definitions of life that are used to interpret observational data, and by the growing number of false positives for traditional biosignatures (Cleland and Chyba, 2002; Benner, 2010; Cleland, 2012; Mix, 2015; Schwieterman et al., 2016; Bich and Green, 2018; Harman and Domagal-Goldman, 2018; Schwieterman et al., 2018; Mariscal and Doolittle, 2020; Janin, 2021; Vickers et al., 2023).
@@
One way in which the community aims to overcome the problems with defining life is to develop “agnostic” biosignatures111While the phrase agnostic biosignatures is relatively new, the concept of designing non-Earth centric life detection experiments dates back to at least 1961 (Sagan, 1961; Lovelock, 1965).—detecting signs of life that are not particular to Earth-life or any other hypothetical instances of life (Marshall et al., 2021; Smith et al., 2021). Yet, agnostic biosignature proposals are sometimes built on restricted concepts of habitability (e.g., requiring rocky, watery planets), or simple anomaly detection (i.e., without using a working definition of life) (Kinney and Kempes, 2022; Cleaves et al., 2023).
@@
A proposed remedy for the issues surrounding traditional biosignatures has been to use so-called “statistical” biosignatures—relying on integrating multiple lines of evidence to increase confidence in a discovery, or generating ensembles of data to better constrain the probability of making specific observations contingent on the presence of life (Lin and Loeb, 2015; Catling et al., 2018; Walker et al., 2018; Affholder et al., 2021; Bixel and Apai, 2021). However, such statistical biosignatures often rely on assumptions about the prior probability of abiogenesis, or on the existence of unambiguous biosignatures for single planets. These approaches are often trying to solve more specific issues, like estimating the frequency of Earth-like life, the frequency of planets originating from panspermia compared to abiogenesis, or finding evidence of a specific metabolic process (Lin and Loeb, 2015; Walker et al., 2018; Affholder et al., 2021; Checlair et al., 2021; Kovacevic, 2022).
@@
Here we ask, can we detect the presence of life if we postulate that life is spreading between and terraforming planets?222Here, “life spreading” refers to interstellar panspermia, and “terraforming planets” refers to modifying observable characteristics of planets. This is an astrobiological “hinge proposition”333An assumption that enables a specific method of inquiry but which the method itself cannot verify. As Kinney and Kempes (2022) write, “…it may still be the case that such assumptions are needed in order for us to have any chance of conducting successful inquiry in astrobiology.” of the kind described by Kinney & Kempes as necessary to make sure “that the door of astrobiology can turn properly” under conditions of deep uncertainty (Kinney and Kempes, 2022). It is also somewhat the converse of a question posed in other panspermia related work: Can the prevalence of panspermia be constrained when assuming we have a way to detect life? (Lin and Loeb, 2015; Balbi and Grimaldi, 2020; Grimaldi et al., 2021; Lingam et al., 2021; Kovacevic, 2022).
@@
In fact, our postulates (of terraforming and panspermia) are less peculiar than they might at first seem: while existing work doesn’t always explicitly refer to terraforming, it must be assumed—in the sense that the mere presence of life stable over geological timescales would create environmental feedback with a planet, and if that life is to be detectable it must modify a planet’s observables. Such planetary modification by life is a well-documented phenomena—e.g., the rise of during the great oxygenation event (Olejarz et al., 2021), or the rise of from human industrial activity (Lynas et al., 2021). Both of these examples represent unintentional terraforming, but intentional terraforming from intelligent life has been an area of active research (Beech et al., 2021; Fogg, 1998; Zubrin and Mckay, 1993), and would be expected to cause environmental change at a faster timescale.
@@
The feasibility of interstellar lithopanspermia (non-intelligent exchange of material via rocks) has been discussed at length in other work (see e.g. Carroll-Nellenback et al. 2019; Grimaldi et al. 2021; Gobat 2021; Totani 2023; Cao et al. 2024 and references therein). Crucially, it appears plausible, although likelihoods and rates of material exchange vary considerably depending on assumptions on timing, amount of material ejected, organismal hardiness, and capture rates, among other features (e.g., recent work estimates that lithopanspermia is a plausible life-seeding mechanism for planets in our Galaxy, but this estimate includes only Earth-sized planets in the habitable zone, Cao et al. 2024). Arguments have likewise been made for the feasibility of directed interstellar panspermia by intelligent life (Wright et al., 2014; Newman and Sagan, 1981; Crick and Orgel, 1973). Ultimately, our postulates of panspermia and terraforming are merely well understood hallmarks of life (proliferation via replication, and adaptation with bi-directional environmental feedback), escalated to the planetary scale, and executed on an interstellar scale.
@@
We use simulations to show that statistical correlations between the spatial distribution of planets around different host stars, and their observable characteristics could itself be evidence of life, without the need for a separate biosignature that could reliably detect life on any given planet in isolation. The agnosticism of our biosignature is inseparable from its emergence at the scale of a population of planets—singleton planetary anomalies might be explained away by unknown geochemical processes, or targeted simply because they are anomalous (without a clear hypothesis of why they should be explained by life). Hypothesizing that life spreads via panspermia and terraforming allows us to search for biosignatures while forgoing any strong assumptions about not only the peculiarities of life (e.g., its metabolism) and planetary habitability (e.g., requiring surface liquid water, Gobat 2021), but even the potential breadth of structure and chemical complexity underpinning living systems (Solé and Munteanu, 2004; Kim et al., 2019; Bartlett et al., 2022; Wong et al., 2023).
@@
We first describe our model (Sec. II.1) and approach for statistically identifying the presence of terraformed planets (Sec. II.2). We show that evolving our model in time can increase the correlation between planets’ compositions and positions (Sec. III.1). To identify specific planets likely to have been terraformed, we cluster planets by observable characteristics, then select clusters which are spatially localized and cause a decrease in correlation when removed (Secs. II.2.3, II.2.4). We evaluate these clusters by calculating how well they correctly classify terraformed planets as being terraformed, and non-terraformed planets as being non-terraformed (Sec. III.2). Finally, we discuss how our results might change due to theoretical and observational constraints, identify specific ways in which better understanding astrophysical and planetary processes could improve life detection, and speculate on the concept of life at the scale of populations of planets.
II Methods
II.1 Modeling Panspermia and Terraforming
II.1.1 Summary
@@
We created an agent-based model to simulate life444In this work, we use life specifically to refer to an agent (used strictly in the “entity” sense of the word agent) traveling from a parent terraformed planet to a target non-terraformed planet, and whose arrival at the target planet triggers terraforming. spreading between planets orbiting different host stars. The model is initialized with 1000 planets (one of which is terraformed/panspermia-capable), fixed in a frame of reference, and uniformly randomly distributed in a continuous 3D volume. All planets are assumed to be in different star systems (i.e., we assume one planet per star). “Planets” have compositions representing the observable characteristics of the planet, and “life” agents have compositions representing the observable characteristics of the planet which the life originates from. During simulation, terraformed planets 555We specifically avoid saying living planets, because our methodology cannot distinguish between living planets incapable of panspermia and non-living planets. emit life towards a target non-terraformed planet at a constant velocity 666This approach is most similar to directed panspermia by intelligent life, but could also be envisioned as undirected lithopanspermia, where we refrain from burdening the simulations with agents that won’t ever be able to interact with planets under the rules of our model.. The target planets are selected based on proximity in position and in composition (see Sec. II.1.4). When life arrives at a planet, the planet’s composition is modified (terraformed) based on both the initial composition of the planet and the composition of the incoming life. Life originating from that terraformed planet will then reflect its new composition. The time that life takes to travel to a planet in our model can be thought of as the sum of the time it takes to travel to and terraform a planet. Terraformed planets emit life at a fixed rate, provided suitable target planets exist (see Sec. II.1.4). The simulation ends when all planets are terraformed, or no suitable target planets remain (Fig. 1B).
II.1.2 Compositions and Terraforming
@@
Both “life” and “planets” have compositions represented by vectors of 10 real numbers (each element ). In our model, these compositions are abstractions, used as a means to compare the compatibility of any given life and planet. While we might in the abstract think of one vector representing the observable characteristics of a planet, this same vector would not represent the same observable characteristics of life compatible with that planet. For life, it instead represents the kind of planetary observable characteristics which are compatible with that life. For example, perhaps a planet’s composition vector corresponds to an atmosphere composed of 95% and 5% . The same composition vector in the context of life is used to refer to a kind of life which could be compatible with a planet with an atmospheric composition of 95% and 5% . This does not mean that life on such a planet is literally composed of 95% and 5% . This captures the idea that if a life and a planet have identical composition vectors, then that life will surely be able to terraform the planet777This is an extreme example where technically life wouldn’t even need to terraform the planet, because it implies that the planet’s observable characteristics already perfectly match the planet from where the life originated.. Correspondingly, in our model a life’s composition is identical to the planet from which it is emitted. Initial compositions of each planet are random, with each element in each composition vector drawn from a continuous uniform distribution . This is a loose approximation of the unknowns surrounding the distribution of diversity of planets between planetary systems (Tasker et al., 2020; Pacetti et al., 2022).
@@
Planet compositions are modified by terraforming, where a mixing function determines how to generate the terraformed composition based on the non-terraformed planet composition and the incoming life’s composition. We apply a mixing function inspired by horizontal gene transfer (Senaratna, 2005; Umbarkar et al., 2015), but other mixing functions could also be used. The most important feature we aim to capture is the bi-directional feedback between the terraforming life, and the terraformed planet—such that future life emitted from the terraformed planet will inherit the post-terraforming characteristics of that planet. We keep randomly chosen element from the planet’s pre-terraformed composition, and the rest of the terraformed composition is inherited from the terraforming life (Fig. 1C).
II.1.3 Model Stepping
@@
Each step of the model increments the time by . Time units are arbitrary, and only impact the distance that life travels in a step, . Here, , and , meaning that for every 10 model steps, a planet can emit 1 new life, and existing life can travel 20 units (for reference, the average distance between planets is units).
II.1.4 Target Planet Selection
@@
The destination selection aims to balance the need of identifying a target planet that is both compositionally similar, and physically near, the parent planet. This is based on the assumption that life directing panspermia would want to minimize travel time and terraforming energy costs. In the context of lithopanspermia, this could be interpreted as an assumption that beyond a certain positional distance capture becomes vanishingly unlikely, and beyond a certain compositional distance compatibility becomes vanishingly unlikely.
@@
The algorithm we used for determining a life’s target planet has a positional distance threshold, . Among all non-terraformed planets, , under that positional distance threshold from the life’s origin planet, it selects the planet nearest in compositional space, . Here, , is shorthand for all non-terraformed planets within of life’s origin planet, and is the planet at minimum compositional distance to the life’s origin planet (Fig. 1A). The distances calculated are Euclidean. A range of positional thresholds, , corresponding to and of the maximum distance possible between any two points in space were chosen for in order to span the transition between a model where terraforming of every planet is impossible and every planet is possible, based on the average distance between planets.
II.2 Identifying the Presence of Terraformed Planets
II.2.1 Summary
@@
We hypothesize that the process of panspermia and terraforming in our model will lead to a population of planets with anomalously high positive correlations between their spatial locations and compositions, compared to random permutations of these planets’ compositions. We quantify this using the Mantel test—a statistical test common in ecological science (Sec. II.2.2). By clustering on the planet compositions, we can identify clusters of planets driving these correlations (Sec. II.2.3). From these initial clusters, we select those localized in space (via their interquartile range, IQR), because we hypothesize that life would not only change the distribution of planetary compositions, but would also do so in a relatively compact portion of the galaxy. We further select clusters which, when removed, cause a decrease in the Mantel coefficient of the residual space of planets (Sec. II.2.4). Finally, we attempt to evaluate how well our clusters reflect the presence of truly terraformed planets (Sec. II.2.5).
II.2.2 Mantel Test